# Cosmic (super)strings with a time-varying tension

### v1

@@ 5 Outlook
In this paper, we analysed the dynamics of a cosmic string network with a time-dependent tension. This is motivated by the idea that, in String Theory constructions, the tensions of both fundamental and composite strings can depend non-trivially on the expectation values of the moduli. In turn, the latter can evolve as a function of time, and induce variations in the string tension. This could be very relevant for the very early stages of the universe’s evolution, which are mostly unconstrained by observations until the onset of Big Bang Nucleosynthesis (BBN) Allahverdi:2020bys and where string theoretic considerations motivate rolling scalars Cicoli:2024bwq . As an example, we discussed the case where a time-dependent volume modulus in the very early universe results in decreasing string scale $m_{s}$ Apers:2024ffe , potentially offering a way to mitigate observational constraints on the existence of cosmic superstrings Conlon:2022pnx .
To this aim, we extended the Velocity One Scale (VOS) model of Martins:1996jp ; Martins:2000cs to describe a string network in the time-dependent case. This was achieved by considering the effect of new terms appearing in the microscopic equations of motion, and through a careful averaging on larger scales. In particular, we provided both microscopic and macroscopic arguments to suggest that the ansatz for the momentum parameter $k(v)$ proposed in Martins:2000cs should also hold in the macroscopic case. The outcome of this analysis was presented in Section 4.2, and we also summarize it here. We found that the well-known (scaling) attractor solution, with a linearly growing correlation length ($L=\xi t$), persists to the case of a varying string tension. Notice, however, that due to a decreasing tension the energy density in the network is no longer a fixed fraction of the background, and the scaling parameters are modified in accordance to Eqs (60)-(61). More surprisingly, we also noticed that if
$2\mathcal{H}+\frac{\dot{\mu}}{\mu}<0,$ (78)
a new attractor solution can exist, corresponding to a network that redshifts like radiation. In that case, the end state of the system depends on the initial conditions, although the standard scaling solution is preferred for realistic values of the latter. In Section 4.3, we also discussed the evolution of the loop population taking gravitational radiation into account, explicitly confirming the finding of Conlon:2024uob on the bizarre phenomena of loop growth when (78) is violated.
@@
  • The VOS model and some of its extensions (such as superconducting strings) have been numerically verified in a number of cases, see for example Moore:2001px ; Hindmarsh:2017qff ; Correia:2019bdl ; Martins:2020jbq . To put our results on firmer grounds, it would be necessary to carry out numerical simulations of cosmic strings with time-varying tensions. This could be used to both test some of our key assumptions (such as the functional form of $k(v)$) and also the final results on the attractor solutions.
@@
  • Our microscopic, analytic understanding of the momentum parameter $k(v)$ is only limited to certain particular configurations, such as the static, infinite string. Is it possible to find a general argument showing that $k(v)$ switches sign when $v^{2}=\frac{1}{2}$, in accordance with the ansatz (33)?
@@
  • It would be interesting to examine more in depth the effect of a varying tension for realistic axion string networks (see Benabou:2023npn for a recent analysis on the cosmology of string theory axion strings). [^21] For completeness, we should mention that a logarithmic violation of the scaling regime has been observed for axion strings Gorghetto:2018myk ; Gorghetto:2020qws , due to the slowly-varying, IR contribution to the tension. Are there any natural string theory realisations of this scenario, and would they lead to any observable signatures?
@@
Moreover, our results open the door to a computation of the gravitational wave spectra originating from time-dependent string networks, which we will report on in the future us:2024? . As already remarked in the introduction, the GW spectrum can only be reliably computed when the evolution of the network is well understood, so our results provide a first step in this direction. If a network of cosmic strings underwent a period in the early Universe on which the tension decreased, as suggested in scenarios featuring volume-modulus kination Conlon:2022pnx ; Conlon:2024ene ; Conlon:2024hgw ; Conlon:2024uob or other displacements of the volume modulus after inflation, as in Burgess:2022nbx ; Cicoli:2024bwq , we expect a modification in the GW spectrum at high frequencies, see Aggarwal:2020olq for a discussion of possible sources and detectors  [^22] Another GW spectrum arising from long fundamental strings, this time in thermal equilibrium (which is possible if the strings feature additional decay channels which prevent the scaling regime) has been studied in Frey:2024jqy .. Modifications in the tilt of the spectrum due to a varying equation of state of the background have been studied before Cui:2017ufi ; Cui:2018rwi ; Gouttenoire:2021jhk ; Servant:2023tua and it would be interesting to compare these predictions with those arising from a period of varying tension.
Acknowledgements.
@@
We are grateful to Anish Ghoshal for participating in the initial stages of this project, as well as collaboration on the related work us:2024? . We would also like to thank Joe Conlon, Ed Copeland and Noelia Sánchez González for conversations on their recent work Conlon:2024uob and Amelia Drew, Joe Conlon, Fernando Quevedo and Paul Shellard for helpful comments on the manuscript and extensive discussions. We also thank the organisers of SUSY 23 and of the workshop WISPs in String Cosmology, where some of this work took place. The research of FR was partially supported by the Dutch Research Council (NWO) via a Start-Up grant and a Vici grant, and by a junior postdoctoral fellowship of the Fonds Wetenschappelijk Onderzoek (FWO), project number 12A1Q25N. The work of GV has been partially supported by STFC consolidated grant ST/P000681/1, ST/T000694/1. GV thanks the Institute for Theoretical Physics of Utrecht University for hospitality. This work is partly based upon work from the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A Dimensional reduction: a primer
One of the main motivations for this article is the fact that the four-dimensional Planck scale is determined dynamically in string theory. In this Appendix, aiming for self-consistency, we review the well-known process of dimensional reduction (see e.g. Rattazzi:2003ea ). Let us consider the simple example of the dimensional reduction of a 5-dimensional massless scalar theory on a circle:
@@
$\left(\partial_{\mu}\partial^{\mu}+\left(\frac{n}{R}\right)^{2}\right)\phi^{(n)}(x)=0\,$ (81)
@@
$S_{4}=2\pi R\int{d^{4}x\,\sum_{n}\left(-\partial_{\mu}\phi^{*\,(n)}\partial^{\mu}\phi^{(n)}-\left(\frac{n}{R}\right)^{2}|\phi^{(n)}|^{2}\right)}\,.$ (82)
@@
$G_{\mu\nu}=\left(\frac{\sigma_{0}}{\sigma}\right)^{1/3}\left(g_{\mu\nu}-\kappa^{2}A_{\mu}A_{\nu}\right)\,,\qquad G_{\mu 5}=-\kappa\sigma^{2/3}A_{\mu}\left(\frac{\sqrt{\alpha^{\prime}}}{R}\right)^{2}\,,\qquad G_{55}=\sigma^{2/3}\left(\frac{\sqrt{\alpha^{\prime}}}{R}\right)^{2}\,.$ (84)
@@
$S_{EH}=\frac{2\pi\sqrt{\alpha^{\prime}}\sigma_{0}^{1/3}\,M_{p,5}^{3}}{2}\int{d^{4}x\,\sqrt{-g}\left(R_{(4)}-\frac{1}{4}\sigma F_{\mu\nu}F^{\mu\nu}-\frac{1}{6}\frac{\partial_{\mu}\sigma\,\partial^{\mu}\sigma}{\sigma^{2}}\right)}+\dots\,,$ (85)
@@ Appendix B Semi-analytic solution for the circular loop
$\varepsilon^{2}=\frac{R^{2}}{1-(\dot{R}L)^{2}}\,,$ (86)
so that $\varepsilon$ coincides with $R$ when the velocity of $R$ is zero, i.e: at its maximum amplitude, and otherwise performs a Lorentz transformation from the center of mass frame of the string so that the proper length is invariant. We thus conclude that varying $\varepsilon$ indicates, in general, that the proper length of the string $L$ changes. For the circular loop, this implies that the maximum value of the function $R(t)$ grows or decreases, on average, and via Eq. (42) it is clear that this depends in the sign of $2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}$.
We can be more concrete and find an analytic expression for the behavior of loops in an arbitrary power-law cosmology. To do so, let us consider the formal solution to Eq. (42):
$\varepsilon(t)=\varepsilon(0)\,e^{-q\int{\frac{d\tau}{\tau}\dot{\textbf{x}}^{2}}}\,$ (87)
@@
$\varepsilon(\tau)=\frac{R(\tau)}{\sqrt{1-(\dot{R}L)^{2}}}\,$ (88)
@@
$\dot{R}(\tau)L=-\sin(f(\tau/L))\,.$ (89)
This implies, via Eq. (87), that the solution for $R(\tau)$ is given implicitly by
$R(\tau)=e^{-q\int{\frac{d\tau}{\tau}\sin(f(\tau/L))^{2}}}\cos(f(\tau/L))\,,$ (90)
@@
    $R_{max}(\tau)\sim\left(\frac{\tau}{\tau_{0}}\right)^{-q/2}\,,$ (91)
@@
We thus find that, regardless of the particular form of $f(\tau/L)$, whether the overall amplitude of the loops increases only depends in the sign of $q$, as pointed out in Conlon:2024uob .
Let us now go further and investigate the behaviour of $f(\tau/L)$. Consistency of equations (89) and (90) requires that $f(\tau/L)$ satisfies
$e^{q\int{\frac{d\tau}{\tau}\left(\sin\left(f(\tau/L)\right)\right)^{2}}}=f^{\prime}(\tau/L)+q\frac{L}{\tau}\cos(f(\tau/L))\sin(f(\tau/L))\,.$ (92)
@@
$R(t)=\frac{\cos(f(\tau/L))}{f^{\prime}(\tau/L)+q\frac{L}{\tau}\cos(f(\tau/L))\sin(f(\tau/L))}\,,$ (93)
@@
$x^{2}f^{\prime\prime}(x)+q\,xf^{\prime}(x)-q\,f(x)=0\,,$ (95)
@@
$2x^{2}f^{\prime\prime}(x)-qxf^{\prime}(x)=0\,,$ (98)
@@ Appendix C Details on the dynamical system
We would like to examine the structure of the autonomous, dynamical system (59), which we rewrite here for ease of presentation:
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=\xi\left[\left(\frac{2}{n}-1\right)+\frac{v^{2}}{2}\left(\frac{4}{n}-m\right)\right]+\frac{\tilde{c}(v)v}{2}\\ \frac{{\rm d}v}{{\rm d}z}=(1-v^{2})\left[\frac{k(v)}{\xi}-\left(\frac{4}{n}-m\right)v\right],\\ \end{cases}$ (101)
@@
$\tilde{c}(v)=\tilde{c}_{\infty}\quad\quad\text{or}\quad\quad\tilde{c}(v)=\tilde{c}(1-v^{2})^{\alpha},\quad 0<\alpha<1.$ (102)
@@
$\xi_{s}=\sqrt{\frac{n}{8}\frac{k(v_{s})\left(k(v_{s})+\tilde{c}\right)}{\left(1-\frac{mn}{4}\right)\left(1-\frac{2}{n}\right)}}\,,\quad\quad\quad\quad v_{s}=\sqrt{\frac{n}{2}\frac{k(v_{s})}{k(v_{s})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{mn}{4}}},$ (103)
which formally corresponds to the usual VOS model attractor. Notice, however, that these are implicit equations, and the existence and number of solutions cruciallly depends on the form of $k(v)$. Assuming the usual ansatz (33), this already tells us that $v_{s}^{2}<1/2$ if $mn<4$, and $11/2<v_{s}^{2}<$ 1 if $mn>4$. More precisely, the possible values of $v_{s}$ can be determined from the zeroes of the function
$\mathcal{F}(v)=k(v)-\frac{\tilde{c}v^{2}}{\frac{2(n-2)}{4-mn}-v^{2}}.$ (104)
@@
$J_{s}=\begin{pmatrix}-\frac{\tilde{c}v_{s}}{2\xi_{s}}&k(v_{s})+\frac{\tilde{c}}{2}\\ -\frac{(1-v_{s}^{2})k(v_{s})}{\xi_{s}^{2}}\,\,\,\,\,\,&(1-v_{s}^{2})\left[\frac{k^{\prime}(v_{s})}{\xi_{s}}-\left(\frac{4}{n}-m\right)\right]\\ \end{pmatrix},$ (105)
where all quantities are understood to be evaluated at the fixed point. For later reference, the trace and determinant of the Jacobian at the point(s) (103) can be expressed as
$\text{Tr}\left[J_{s}\right]=-\frac{\tilde{c}v_{s}}{2\xi_{s}}+(1-v_{s}^{2})\left[\frac{k^{\prime}(v_{s})}{\xi_{s}}-\left(\frac{4}{n}-m\right)\right],$ (106)
@@
$\text{Det}\left[J_{s}\right]=\frac{\left(1-v_{s}^{2}\right)\left[4\tilde{c}v_{s}(n-2)(4-mn)-k^{\prime}(v_{s})\left(v_{s}^{2}(4-mn)-2(n-2)\right)^{2}\right]}{2\tilde{c}n^{2}v_{s}}.$ (107)
respectively, where in the second equation we eliminated $\xi_{s}$ for convenience (through the vanishing of the RHS in (59)).
Let begin by discussing the case $mn<4$. Since $\mathcal{F}(0)>0,\mathcal{F}(\sqrt{2}/2)<0$ and $\mathcal{F}^{\prime}(v)<0$ for $20\leq v^{2}\leq 1/$ 2, there always exists exactly one solution in the correct range. [^23] We are also assuming $n>2$. If $k^{\prime}(v_{s})<0$, as is true for any fixed point with $mn<4$, (106) implies $\text{Tr}\left[J_{s}\right]<0$, and from (107) $\text{Det}\left[J_{s}\right]>0$. Therefore, this fixed point is always stable whenever it exists, matching with the expectations for the constant tension case, $m=0$.
For $mn>4$, however, there are always two solutions to $\mathcal{F}(v)=0$ , as $\mathcal{F}(\sqrt{2}/2),\mathcal{F}(1)>0$, and $\mathcal{F}(v)$ has a single minimum in the interval $11/2<v^{2}<$ 1. The one closer to $v^{2}=1/2$ has $k^{\prime}(v)<0$, while the one closer to $v^{2}=1$ has $k^{\prime}(v)>0$. If $k^{\prime}(v_{s})>0$, $\text{Det}\left[J_{s}\right]<0$, and the latter fixed point can never be stable. For the remaining one to be stable, two conditions need to be satisfied: $\text{Det}\left[J_{s}\right]>0$, and $\text{Tr}\left[J_{s}\right]<0$. Using the form of $\mathcal{F}(v)$ given by (104), these translate to [^24] Assuming $k(v_{s})<0$.
$\frac{k^{\prime}(v_{s})}{k^{2}(v_{s})}<-\frac{4(n-2)}{mn-4}\frac{1}{\tilde{c}v_{s}^{3}}$ (108)
@@
$2(1-v_{s}^{2})k(v_{s})>v_{s}\left[2(1-v_{s}^{2})k^{\prime}(v_{s})-\tilde{c}v_{s}\right].$ (109)
respectively. It can be verified numerically, that for reasonable values of $\tilde{c}$ and for $mn>4$, the above conditions are always satisfied with the form of $k(v)$ given in (33). Therefore, the scaling fixed point with the lowest value of $v_{s}^{2}$ is always stable.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=\frac{\tilde{c}_{\infty}n}{2(n-2)n+mn-4}.$ (110)
@@
$J_{r}=\begin{pmatrix}\frac{4}{n}-\frac{m}{2}-1&\xi_{r}\left(\frac{4}{n}-m\right)+\frac{\tilde{c}_{\infty}+\tilde{c}^{\prime}(1)}{2}\\ 0&2\left(\frac{4}{n}-m\right)\\ \end{pmatrix},$ (111)
@@
$mn>{\rm Max}\left\{4,4+n\left(\frac{m}{2}-1\right)\right\}.$ (112)
@@
$\xi=\frac{\tilde{c}_{\infty}n}{2(n-2n+mn-4}+\left[\xi_{0}-\frac{\tilde{c}_{\infty}n}{2(n-2)+mn-4}\right]\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]},$ (113)

### v2

@@ 5 Outlook
In this paper, we analysed the dynamics of a cosmic string network with a time-dependent tension. This is motivated by the idea that, in String Theory constructions, the tensions of both fundamental and composite strings can depend non-trivially on the expectation values of the moduli. In turn, the latter can evolve as a function of time, and induce variations in the string tension. This could be very relevant for the very early stages of the universe’s evolution, which are mostly unconstrained by observations until the onset of Big Bang Nucleosynthesis (BBN) [101] and where string theoretic considerations motivate rolling scalars [52]. As an example, we discussed the case where a time-dependent volume modulus in the very early universe results in decreasing string scale $m_{s}$ [66], potentially offering a way to mitigate observational constraints on the existence of cosmic superstrings [62].
To this aim, we extended the Velocity One Scale (VOS) model of [36, 37] to describe a string network in the time-dependent case. This was achieved by considering the effect of new terms appearing in the microscopic equations of motion, and through a careful averaging on larger scales. In particular, we provided both microscopic and macroscopic arguments to suggest that the ansatz for the momentum parameter $k(v)$ proposed in [37] should also hold in the macroscopic case. The outcome of this analysis was presented in Section 4.2, and we also summarize it here. We found that the well-known (scaling) attractor solution, with a linearly growing correlation length ($L=\xi t$), persists to the case of a varying string tension. Notice, however, that due to a decreasing tension the energy density in the network is no longer a fixed fraction of the background, and the scaling parameters are modified in accordance to Eqs (4.22)-(4.23). More surprisingly, we also noticed that if
$2\mathcal{H}+\frac{\dot{\mu}}{\mu}<0,$ (5.1)
a new attractor solution can exist, corresponding to a network that redshifts like radiation. In that case, the end state of the system depends on the initial conditions, although the standard scaling solution is preferred for realistic values of the latter. In Section 4.3, we also discussed the evolution of the loop population taking gravitational radiation into account, explicitly confirming the finding of [35] on the bizarre phenomena of loop growth when (5.1) is violated.
@@
  • The VOS model and some of its extensions (such as superconducting strings) have been numerically verified in a number of cases, see for example [102, 103, 104, 105, 106]. To put our results on firmer grounds, it would be necessary to carry out numerical simulations of cosmic strings with time-varying tensions. This could be used to both test some of our key assumptions (such as the functional form of $k(v)$) and also the final results on the attractor solutions.
@@
  • Our microscopic, analytic understanding of the momentum parameter $k(v)$ is only limited to certain particular configurations, such as the static, infinite string. Is it possible to find a general argument showing that $k(v)$ switches sign when $v^{2}=\frac{1}{2}$, in accordance with the ansatz (3.14)?
@@
  • It would be interesting to examine more in depth the effect of a varying tension for realistic axion string networks (see [87] for a recent analysis on the cosmology of string theory axion strings). [^23] For completeness, we should mention that a logarithmic violation of the scaling regime has been observed for axion strings [11, 12], due to the slowly-varying, IR contribution to the tension. Are there any natural string theory realisations of this scenario, and would they lead to any observable signatures?
Moreover, our results open the door to a computation of the gravitational wave spectra originating from time-dependent string networks, which we will report on in the future [38]. On a qualitative level, we expect this scenario to release a larger signal at high frequencies, when compared to the constant tension case. The reasoning is as follows: fixing a late-time tension $G\mu_{0}$, the low-frequency spectra of our scenario and a standard scenario with constant tension $G\mu_{0}$ should be indistinguishable, since low frequencies are sensitive to late times. However, a key feature of GWs from cosmic strings is that they can carry information about the full cosmic history, including the early times at high frequencies. At high frequencies, thus, it is natural to expect a growth of the signal as compared with the constant tension case, since GW spectra arising from networks of higher tension are larger. The spectrum, however, depends on the evolution of the loop distribution, which is qualitatively different when the tension varies, so further work is required.
@@
As already remarked in the introduction, the GW spectrum can only be reliably computed when the evolution of the network is well understood, so our results provide a first step in this direction. If a network of cosmic strings underwent a period in the early Universe on which the tension decreased, as suggested in scenarios featuring volume-modulus kination [62, 68, 67, 35] or other displacements of the volume modulus after inflation, as in [51, 52], we expect a modification in the GW spectrum at high frequencies, see [107] for a discussion of possible sources and detectors. [^24] Another GW spectrum arising from long fundamental strings, this time in thermal equilibrium (which is possible if the strings feature additional decay channels which prevent the scaling regime) has been studied in [108]. Modifications in the tilt of the spectrum due to a varying equation of state of the background have been studied before [25, 26, 109, 110, 111] and it would be interesting to compare these predictions with those arising from a period of varying tension.
Acknowledgments
@@
We are grateful to Anish Ghoshal for participating in the initial stages of this project, as well as collaboration on the related work [38]. We would also like to thank Joe Conlon, Ed Copeland and Noelia Sánchez González for conversations on their recent work [35] and Amelia Drew, Joe Conlon, Fernando Quevedo, Paul Shellard and the anonymous referee for helpful comments on the manuscript and extensive discussions. We also thank the organisers of SUSY 23 and of the workshop WISPs in String Cosmology, where some of this work took place. The research of FR was partially supported by the Dutch Research Council (NWO) via a Start-Up grant and a Vici grant, and by a junior postdoctoral fellowship of the Fonds Wetenschappelijk Onderzoek (FWO), project number 12A1Q25N. The work of GV has been partially supported by STFC consolidated grant ST/P000681/1, ST/T000694/1. GV thanks the Institute for Theoretical Physics of Utrecht University for hospitality. This work is partly based upon work from the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A Dimensional reduction: a primer
One of the main motivations for this article is the fact that the four-dimensional Planck scale is determined dynamically in string theory. In this Appendix, aiming for self-consistency, we review the well-known process of dimensional reduction (see e.g. [112]). Let us consider the simple example of the dimensional reduction of a 5-dimensional massless scalar theory on a circle:
@@
$\left(\partial_{\mu}\partial^{\mu}+\left(\frac{n}{R}\right)^{2}\right)\phi^{(n)}(x)=0\,$ (A.3)
@@
$S_{4}=2\pi R\int{d^{4}x\,\sum_{n}\left(-\partial_{\mu}\phi^{*\,(n)}\partial^{\mu}\phi^{(n)}-\left(\frac{n}{R}\right)^{2}|\phi^{(n)}|^{2}\right)}\,.$ (A.4)
@@
$G_{\mu\nu}=\left(\frac{\sigma_{0}}{\sigma}\right)^{1/3}\left(g_{\mu\nu}-\kappa^{2}A_{\mu}A_{\nu}\right)\,,\qquad G_{\mu 5}=-\kappa\sigma^{2/3}A_{\mu}\left(\frac{\sqrt{\alpha^{\prime}}}{R}\right)^{2}\,,\qquad G_{55}=\sigma^{2/3}\left(\frac{\sqrt{\alpha^{\prime}}}{R}\right)^{2}\,.$ (A.6)
@@
$S_{EH}=\frac{2\pi\sqrt{\alpha^{\prime}}\sigma_{0}^{1/3}\,M_{p,5}^{3}}{2}\int{d^{4}x\,\sqrt{-g}\left(R_{(4)}-\frac{1}{4}\sigma F_{\mu\nu}F^{\mu\nu}-\frac{1}{6}\frac{\partial_{\mu}\sigma\,\partial^{\mu}\sigma}{\sigma^{2}}\right)}+\dots\,,$ (A.7)
@@ Appendix B Semi-analytic solution for the circular loop
$\varepsilon^{2}=\frac{R^{2}}{1-(\dot{R}L)^{2}}\,,$ (B.1)
so that $\varepsilon$ coincides with $R$ when the velocity of $R$ is zero, i.e: at its maximum amplitude, and otherwise performs a Lorentz transformation from the center of mass frame of the string so that the proper length is invariant. We thus conclude that varying $\varepsilon$ indicates, in general, that the proper length of the string $L$ changes. For the circular loop, this implies that the maximum value of the function $R(t)$ grows or decreases, on average, and via Eq. (4.4) it is clear that this depends in the sign of $2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}$.
We can be more concrete and find an analytic expression for the behavior of loops in an arbitrary power-law cosmology. To do so, let us consider the formal solution to Eq. (4.4):
$\varepsilon(t)=\varepsilon(0)\,e^{-q\int{\frac{d\tau}{\tau}\dot{\textbf{x}}^{2}}}\,$ (B.2)
@@
$\varepsilon(\tau)=\frac{R(\tau)}{\sqrt{1-(\dot{R}L)^{2}}}\,$ (B.3)
@@
$\dot{R}(\tau)L=-\sin(f(\tau/L))\,.$ (B.4)
This implies, via Eq. (B.2), that the solution for $R(\tau)$ is given implicitly by
$R(\tau)=e^{-q\int{\frac{d\tau}{\tau}\sin(f(\tau/L))^{2}}}\cos(f(\tau/L))\,,$ (B.5)
@@
    $R_{max}(\tau)\sim\left(\frac{\tau}{\tau_{0}}\right)^{-q/2}\,,$ (B.6)
@@
We thus find that, regardless of the particular form of $f(\tau/L)$, whether the overall amplitude of the loops increases only depends in the sign of $q$, as pointed out in [35].
Let us now go further and investigate the behaviour of $f(\tau/L)$. Consistency of equations (B.4) and (B.5) requires that $f(\tau/L)$ satisfies
$e^{q\int{\frac{d\tau}{\tau}\left(\sin\left(f(\tau/L)\right)\right)^{2}}}=f^{\prime}(\tau/L)+q\frac{L}{\tau}\cos(f(\tau/L))\sin(f(\tau/L))\,.$ (B.7)
@@
$R(t)=\frac{\cos(f(\tau/L))}{f^{\prime}(\tau/L)+q\frac{L}{\tau}\cos(f(\tau/L))\sin(f(\tau/L))}\,,$ (B.8)
@@
$x^{2}f^{\prime\prime}(x)+q\,xf^{\prime}(x)-q\,f(x)=0\,,$ (B.10)
@@
$2x^{2}f^{\prime\prime}(x)-qxf^{\prime}(x)=0\,,$ (B.13)
@@ Appendix C Details on the dynamical system
We would like to examine the structure of the autonomous, dynamical system (4.21), which we rewrite here for ease of presentation:
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=\xi\left[\left(\frac{2}{n}-1\right)+\frac{v^{2}}{2}\left(\frac{4}{n}-m\right)\right]+\frac{\tilde{c}(v)v}{2}\\ \frac{{\rm d}v}{{\rm d}z}=(1-v^{2})\left[\frac{k(v)}{\xi}-\left(\frac{4}{n}-m\right)v\right],\\ \end{cases}$ (C.1)
@@
$\tilde{c}(v)=\tilde{c}_{\infty}\quad\quad\text{or}\quad\quad\tilde{c}(v)=\tilde{c}(1-v^{2})^{\alpha},\quad 0<\alpha<1.$ (C.2)
@@
$\xi_{s}=\sqrt{\frac{n}{8}\frac{k(v_{s})\left(k(v_{s})+\tilde{c}\right)}{\left(1-\frac{mn}{4}\right)\left(1-\frac{2}{n}\right)}}\,,\quad\quad\quad\quad v_{s}=\sqrt{\frac{n}{2}\frac{k(v_{s})}{k(v_{s})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{mn}{4}}},$ (C.3)
which formally corresponds to the usual VOS model attractor. Notice, however, that these are implicit equations, and the existence and number of solutions cruciallly depends on the form of $k(v)$. Assuming the usual ansatz (3.14), this already tells us that $v_{s}^{2}<1/2$ if $mn<4$, and $11/2<v_{s}^{2}<$ 1 if $mn>4$. More precisely, the possible values of $v_{s}$ can be determined from the zeroes of the function
$\mathcal{F}(v)=k(v)-\frac{\tilde{c}v^{2}}{\frac{2(n-2)}{4-mn}-v^{2}}.$ (C.4)
@@
$J_{s}=\begin{pmatrix}-\frac{\tilde{c}v_{s}}{2\xi_{s}}&k(v_{s})+\frac{\tilde{c}}{2}\\ -\frac{(1-v_{s}^{2})k(v_{s})}{\xi_{s}^{2}}\,\,\,\,\,\,&(1-v_{s}^{2})\left[\frac{k^{\prime}(v_{s})}{\xi_{s}}-\left(\frac{4}{n}-m\right)\right]\\ \end{pmatrix},$ (C.5)
where all quantities are understood to be evaluated at the fixed point. For later reference, the trace and determinant of the Jacobian at the point(s) (C.3) can be expressed as
$\text{Tr}\left[J_{s}\right]=-\frac{\tilde{c}v_{s}}{2\xi_{s}}+(1-v_{s}^{2})\left[\frac{k^{\prime}(v_{s})}{\xi_{s}}-\left(\frac{4}{n}-m\right)\right],$ (C.6)
@@
$\text{Det}\left[J_{s}\right]=\frac{\left(1-v_{s}^{2}\right)\left[4\tilde{c}v_{s}(n-2)(4-mn)-k^{\prime}(v_{s})\left(v_{s}^{2}(4-mn)-2(n-2)\right)^{2}\right]}{2\tilde{c}n^{2}v_{s}}.$ (C.7)
respectively, where in the second equation we eliminated $\xi_{s}$ for convenience (through the vanishing of the RHS in (4.21)).
Let begin by discussing the case $mn<4$. Since $\mathcal{F}(0)>0,\mathcal{F}(\sqrt{2}/2)<0$ and $\mathcal{F}^{\prime}(v)<0$ for $20\leq v^{2}\leq 1/$ 2, there always exists exactly one solution in the correct range. [^25] We are also assuming $n>2$. If $k^{\prime}(v_{s})<0$, as is true for any fixed point with $mn<4$, (C.6) implies $\text{Tr}\left[J_{s}\right]<0$, and from (C.7) $\text{Det}\left[J_{s}\right]>0$. Therefore, this fixed point is always stable whenever it exists, matching with the expectations for the constant tension case, $m=0$.
For $mn>4$, however, there are always two solutions to $\mathcal{F}(v)=0$ , as $\mathcal{F}(\sqrt{2}/2),\mathcal{F}(1)>0$, and $\mathcal{F}(v)$ has a single minimum in the interval $11/2<v^{2}<$ 1. The one closer to $v^{2}=1/2$ has $k^{\prime}(v)<0$, while the one closer to $v^{2}=1$ has $k^{\prime}(v)>0$. If $k^{\prime}(v_{s})>0$, $\text{Det}\left[J_{s}\right]<0$, and the latter fixed point can never be stable. For the remaining one to be stable, two conditions need to be satisfied: $\text{Det}\left[J_{s}\right]>0$, and $\text{Tr}\left[J_{s}\right]<0$. Using the form of $\mathcal{F}(v)$ given by (C.4), these translate to [^26] Assuming $k(v_{s})<0$.
$\frac{k^{\prime}(v_{s})}{k^{2}(v_{s})}<-\frac{4(n-2)}{mn-4}\frac{1}{\tilde{c}v_{s}^{3}}$ (C.8)
@@
$2(1-v_{s}^{2})k(v_{s})>v_{s}\left[2(1-v_{s}^{2})k^{\prime}(v_{s})-\tilde{c}v_{s}\right].$ (C.9)
respectively. It can be verified numerically, that for reasonable values of $\tilde{c}$ and for $mn>4$, the above conditions are always satisfied with the form of $k(v)$ given in (3.14). Therefore, the scaling fixed point with the lowest value of $v_{s}^{2}$ is always stable.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=\frac{\tilde{c}_{\infty}n}{2(n-2)n+mn-4}.$ (C.10)
@@
$J_{r}=\begin{pmatrix}\frac{4}{n}-\frac{m}{2}-1&\xi_{r}\left(\frac{4}{n}-m\right)+\frac{\tilde{c}_{\infty}+\tilde{c}^{\prime}(1)}{2}\\ 0&2\left(\frac{4}{n}-m\right)\\ \end{pmatrix},$ (C.11)
@@
$mn>{\rm Max}\left\{4,4+n\left(\frac{m}{2}-1\right)\right\}.$ (C.12)
@@
$\xi=\frac{\tilde{c}_{\infty}n}{2(n-2n+mn-4}+\left[\xi_{0}-\frac{\tilde{c}_{\infty}n}{2(n-2)+mn-4}\right]\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]},$ (C.13)